apply method of variation of parameters to solve D.E.
a)x^2y''+xy'-y=x^2e^x
1
Expert's answer
2020-03-24T13:06:05-0400

This equation is euler-cauchy equation.

Substitute "y=x^{\lambda}" into the homogenous equation

"x^2(x^\lambda)''+x(x^\lambda)'-x^\lambda=0"

"\lambda(\lambda-1)x^\lambda+\lambda x^\lambda- x^\lambda=0"

"\lambda^2-1=0; \lambda=\pm1"

Hence the solution of the homogenous equation

"y=c_1x+c_2x^{-1}"

Using method of variations of parameters "c_1=c_1(x);c_2=c_2(x)" will give us the system

"\begin{matrix}
 c_1'x+c_2'x^{-1}=0 \\
 c_1'-c_2'x^{-2}=e^x 
\end{matrix}\implies
\begin{matrix}
 c_1'=c_2'x^{-2}+e^x \\
 xe^x+c_2'x*x^{-2}+c_2'x^{-1}=0 
\end{matrix}"

"c_2'x^{-1}=-xe^x\implies c_2'=-\frac{1}{2}x^2e^x"

"c_2=-\frac{1}{2}e^x(x^2-2x+2)"

"c_1'=-\frac{1}{2}x^2e^x*x^{-2}+e^x=\frac{1}{2}e^x"

"c_1=\frac{1}{2}e^x"

Therefore

"y=c_1x+c_2x^{-1}+\frac{1}{2}xe^x-\frac{1}{2}x^{-1}e^x(x^2-2x+2)"